40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu May 16 04:46:37 2013, Jenne, Update, LSC, Kiwamu's sensing matrix measurement script revived 
    Reply  Thu May 16 23:48:39 2013, Jenne, Update, LSC, Kiwamu's sensing matrix measurement script revived 
       Reply  Mon May 20 17:49:36 2013, Jenne, Update, LSC, PRMI sensing matrix - not high quality data PRMIsensMatQuickAnalysis.m.gz
          Reply  Mon May 20 18:47:47 2013, Koji, Update, LSC, PRMI sensing matrix - not high quality data 
             Reply  Tue May 21 14:48:08 2013, Jenne, Update, LSC, PRMI sensing matrix - not high quality data 
       Reply  Mon May 20 18:50:22 2013, Jenne, Update, LSC, Kiwamu's sensing matrix measurement script revived 
          Reply  Tue May 21 18:22:18 2013, Jenne, Update, LSC, Sensing matrix scripts modified to include actuator calibration 
             Reply  Wed May 22 00:08:19 2013, Koji, Update, LSC, Sensing matrix scripts modified to include actuator calibration 
                Reply  Wed May 22 18:24:19 2013, Jenne, Update, SUS, Violin mode survey 
                Reply  Wed May 22 20:50:26 2013, Jenne, Update, LSC, Sensing matrix scripts don't calculate correctly 
                   Reply  Thu May 23 14:45:08 2013, Jenne, Update, LSC, Sensing matrix scripts calculations make more sense now SensMat_JenneMeas_23May2013.pngSensMat_KojiMeas_23May2013.png
                      Reply  Thu May 23 21:45:51 2013, Jenne, Update, LSC, Sensing matrix scripts now check for lockloss 
                Reply  Thu May 23 19:09:15 2013, Jenne, Update, LSC, Sensing matrix scripts modified to include actuator calibration SensMat_21May2013.png
          Reply  Wed May 22 18:07:36 2013, Jenne, Update, LSC, Kiwamu's sensing matrix measurement script revived 
             Reply  Thu May 23 00:49:13 2013, Jenne, Update, LSC, LSC filters loaded 
Message ID: 8630     Entry time: Thu May 23 14:45:08 2013     In reply to: 8621     Reply to this: 8635
Author: Jenne 
Type: Update 
Category: LSC 
Subject: Sensing matrix scripts calculations make more sense now 

I think I have most of the magnitude issues figured out now. 

First of all, the lockin outputs are different from the actual responses in the PDs by a factor of 2. 

If the optic is driven with amplitude D, it will have a response of Asin(wt) + Bcos(wt) + other frequency junk.  The lockin bandpasses the response to get rid of the 'other frequency junk'.  Then creates 2 new signals, one multiplied by cos(wt), the other multiplied by sin(wt).  So, now we have Asin^2(wt) + Bcos(wt)sin(wt) and Asin(wt)cos(wt) + Bcos^2(wt).   If I rewrite these, I have A/2*(1-cos(2wt))+B/2*(sin(2wt) and A/2*sin(2wt)+B/2*(1+cos(2wt)).  We lowpass to get rid of the 2w components, and are left with A/2 for the Q-phase, and B/2 for the I-phase of the lockin outputs.  Since the real amplitudes of the response were A for the Q-phase and B for the I-phase, we need to multiply the lockin outputs by 2.

The other problem was that in the 'uncalibrated' version of numbers that I was printing to compare with Koji's, I had not normalized by the drive amplitude yet.  That happens in the "calibration" part of my script.  So, if I go back to comparing the calibrated versions of our numbers, I get quite close to Koji's answers.

For the PRCL magnitudes, 3 of the 4 numbers match to ~5%.  However, the MICH magnitudes all seem to be off by a factor of 2.  I'm still stuck on this factor of 2, but I'm thinking about it. Also, the phases that Koji and I get are pretty different.

Koji's sensing matrix:

Sensing Matrix, units = cts/meter, phase in degrees
            PRCL Mag   PRCL Phase    MICH Mag   MICH Phase  
REFL33_I    3.100E+11  -109.727      4.900E+09    74.434     
REFL33_Q    3.300E+09  -141.730      7.900E+08    71.120     
REFL55_I    3.200E+11  -109.672      5.900E+08    77.313     
REFL55_Q    1.200E+10  -143.169      6.500E+09    91.559  

My sensing matrix:

Sensing Matrix, units = cts/meter, phase in degrees
            PRCL Mag   PRCL Phase    MICH Mag   MICH Phase  
REFL33_I    3.242E+11  -157.846      1.067E+10    19.010     
REFL33_Q    2.217E+09  -104.088      1.683E+09    28.731     
REFL55_I    3.371E+11  -157.915      3.645E+09    11.072     
REFL55_Q    1.213E+10  -118.348      1.346E+10     2.847

Here are the plotted versions of these matricies:

SensMat_KojiMeas_23May2013.png

SensMat_JenneMeas_23May2013.png

SOME EDITS:  Koji's measurement was 1Hz away from the violin mode, while mine (him running my script) was at the violin mode, so the sensor TFs were actually taken at slightly different frequencies. This helps explain the discrepancies.

Also, the phase in these plots isn't correct, so I need to figure that out. Corrected version of the 'koji' measurement put in place of the incorrect one.  I convert from radians to degrees for my script, but Koji had already reported his phases in degrees, so when I multiplied by 180/pi, it didn't make any sense. I now convert his numbers to radians before running them through my analysis script.

 

ELOG V3.1.3-